A 1500-W heater is designed to be plugged into a 120-V outlet.
What current will flow through the heating coil when the heater is plugged in? I = A
What is R , the resistance of the heater?
How long does it take to raise the temperature of the air in a good-sized living room 3.0m x 5.0m x 8m by 10 degrees C? Note that the specific heat of air is 1006 J/(kg x degrees C) and the density of air is 1.20 kg/m^3.
Answer #1
Power = I*V
Solve for I

R = V/I

Power * Time = Mass * (specific heat) * (change in T)

Solve for time.

The answer you get will be an underestimate because heat is also being lost to the walls and windows of the room.

Show your work if you need further assistance
answered by: PRAMIT